In a recent election in Texas for a seat on the United States House of Representatives, Max Sandlin received 25,557 more votes than opponent Dennis Boerner. If the total number of votes was 135,821, find the number of votes for each candidate.
I need to turn it into an equation and solve.
thanks

Found 2 solutions by checkley71, bucky:
Answer by checkley71(8403)

X+(X+25,557)=135,821
2X+25,557=135,821
2X=135,821-25,557
2X=110,264
X=110,264/2
X=55,132 NUMBER OF VOTES FOR THE LOOSER.
55,132+25,557=80,689
55,132+80,689=135,821
135,821=135,821

Answer by bucky(2100)

Let S represent the number of votes that Max Sandlin got.
.
And let B represent the number of votes that Dennis Boerner got.
.
The combined number of votes is 135,821. So the sum of the Sandlin votes and the Boerner votes
is 135821. In equation form this is:
.
S + B = 135821
.
The problem further states that the difference between the Sandlin votes and the Boerner votes
was 25557 with Sandlin getting more votes. In equation form this is:
.
S - B = 25557
.
You can solve this equation for S by adding B to both sides of the equation to get:
.
S = 25557 + B
.
Substitute this value (the right side) for S in the equation for the total votes to change
the total vote equation to:
.
(25557 + B) + B = 135821
.
Add the two B terms on the left side to get:
.
25557 + 2B = 135821
.
Get rid of the 25557 on the left side by subtracting 25557 from both sides to reduce the
equation to:
.
2B = 110264
.
Solve for B by dividing both sides of this equation by 2 to get:
.
B = 110264/2 = 55132
.
So Dennis Boerner got 55132 of the total 135821 votes. This means that Max Sandlin got
the remainder. So Max Sandlin got 135821 - 55132 votes. This means that Sandlin got
80689 votes.
.
Hope this is what you were looking for in the way of equations that enable you to solve the
problem.
.